OpenStudy (anonymous):

Consider the equation 7x + 3y = 42. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. I only need help with part 2! please and thank you. please put it in words

OpenStudy (saifoo.khan):

Part 2. 7x + 3y = 42. Whenever we have to graph it, the easy method is: When u have to find the x-intercept, make y = 0 When u have to find the y-intercept, make x = 0 Dost that makes sense? I will elaborate if u want..

OpenStudy (anonymous):

it does somewhat...

OpenStudy (anonymous):

find both x and y. so 7(0)+3y=42 3y=42 then you divide by 3. you get y=14. Find the x. 7x+3(0)=42 7x=42 then you divide by 7. you get x=6. so x=6 and y=14 is my final answer.

OpenStudy (saifoo.khan):

There^

OpenStudy (anonymous):

Thank You both :)

OpenStudy (saifoo.khan):

Welcome.